OpenStudy (anonymous):

Rachel works part-time from home as a data entry professional. She enters data in forms and is paid based on the entry accuracy rate. The table below shows her compensation. Based on her regular performance, the equation below can be used to determine C, the total amount Rachel earns a month entering data, where x represents the total number of forms she completes that month. C = 17(0.9x) + 13(0.05x) + 0(0.05x) What is the minimum number of forms she must complete in a month to earn a total of at least $1,276? 59 70 79 80

OpenStudy (anonymous):

C = 17(0.9x) + 13(0.05x) + 0(0.05x) Your equation has a 0 in the " + 0(0.05x)" are u sure u typed it correctly?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay i continue to solve it then, C = 17(0.9x) + 13(0.05x) + 0(0.05x) C= 15.3x+0.065x +0 c= 15.365x C minimum 1276 dollars. 1276=15.365x x=83.045 (round up) = 84 forms. Which is none of the answer there, so i beg u to please check the question. there is something wrong.